Simpson drives his car with an average velocity of 48.0 km/hto the east. How long will it take hime to drive 144 km on astraight highway? how much time would he save by increasing hiaverage velocity to 56.0 km/h to the east?
1
Expert's answer
2012-09-18T12:12:14-0400
It will take

T1 = 144/48 = 3 hours

to drive 144 km on a straight highway.

He will save

T = T1 - 144/56 = 3 - 144/56 = 3/7 hour

if he'll increase his average velocity to 56.0 km/h.